# DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing

### v2

@@ 5.6. User Adoption and Impact
6. Discussion
Reflecting on our evaluation results, we observe significant variations in output quality between different candidate plans, underscoring the promise of agentic approaches in optimizing LLM-powered data processing pipelines. However, these findings also highlight the challenges inherent in optimization for such systems. The optimization process in DocETL is fundamentally a best-effort approach, relying on AI agents that, while powerful, are not infallible. The space of possible plan interpretations and decompositions is vast—potentially infinite—and we are limited by the agents’ ability to navigate this space effectively. For instance, when applying a projection synthesis rule (Section 3.3), an agent could conceivably generate any number of map operations to augment or focus the context for the main operation. The quality of the resulting pipeline thus heavily depends on the agent’s ability to guide good decompositions and craft effective prompts for such decompositions.
This reliance on AI agents for optimization suggests the need for a more human-in-the-loop approach. Human intuition can be invaluable in creating prompts for subtasks, or in identifying promising decomposition strategies. This is particularly important given the unintuitive nature of the Pareto frontier of plans, as observed in our results and noted in prior work (Liu et al., 2024b). The fact that some plans are both cheaper and better in terms of output quality challenges conventional optimization heuristics and makes it difficult to devise effective strategies for searching the space of possible plans.
Another difficult challenge is that people often face uncertainty when defining their requirements for AI-powered pipelines. Their needs may only become clear after seeing initial outputs from the system (Shankar et al., 2024b, a). This calls for a more interactive approach to pipeline building, where users can iteratively refine their pipelines based on intermediate results. Additionally, allowing users to influence or even write the validator prompts themselves could provide a powerful mechanism to incorporate domain knowledge and specific evaluation criteria into the optimization process.
It’s important to note that DocETL is not intended for tasks requiring the synthesis of new knowledge beyond what is contained in the input data or provided in operation prompts. Its strength lies in extracting, transforming, and analyzing existing information within large, unstructured datasets—–tasks that traditionally might require significant manual effort and teams of human annotators or domain experts (Nigatu et al., 2023). This allows DocETL to excel in scenarios like our police misconduct analysis case, where it can process vast amounts of unstructured data to reveal patterns and insights.
As we continue to develop DocETL, we are addressing several engineering challenges to enhance its functionality and accessibility. These include implementing robust provenance tracking to allow users to trace analyses back to source documents, enabling compatibility with local LLMs for increased privacy and reduced cloud dependency, and improving scalability for larger datasets.
@@ 7. Related Work
LLM-powered data processing frameworks have recently gained significant attention in the database community. LOTUS (Patel et al., 2024) extends Pandas with semantic operators, while Palimpzest (Liu et al., 2024b) provides a declarative framework focusing on map-like operations. Aryn (Anderson et al., 2024) offers a Spark-like API with PDF extraction capabilities and human-in-the-loop processing. These systems employ various cost-based optimizations, including classical techniques like predicate pushdown (Hellerstein and Stonebraker, 2005) and ML-specific approaches like model cascades (Wang et al., 2017). However, they assume that capable LLMs will produce sufficiently accurate results for user-defined operations. In practice, even state-of-the-art models fall short on complex document processing tasks. DocETL addresses this limitation through agent-driven optimization, exploring decomposition to improve accuracy. Moreover, DocETL is, uniquely, the only system to support documents with lengths that exceed LLM context windows.
Other LLM-powered data processing systems focus on different settings. ZenDB (Lin et al., 2024) optimizes SQL queries for templatized documents, while DocETL handles arbitrary document formats. EVAPORATE (Arora et al., 2023) specializes in table extraction through code synthesis (only where applicable, in semi-structured settings), which could complement DocETL. Regarding LLM agents: Caesura (Urban and Binnig, 2024) uses LLMs to translate natural language to SQL pipelines but leaves optimization for future work; CleanAgent (Qi and Wang, 2024) uses agents to standardize and clean data (and also does not consider optimization). Other systems propose specialized pipelines for specific tasks: for instance, Edge et al. (2024) use a fixed map-reduce pipeline with predefined prompts for knowledge graph querying—whereas DocETL enables flexible pipeline construction and optimization for any document processing task. While prompt optimization (Wen et al., 2024; Khattab et al., 2024) could also complement DocETL, it falls short on complex document tasks, even with human guidance (White et al., 2023), particularly when data and tasks exceed single LLM call capabilities.
LLMs have been leveraged for a variety of data tasks beyond document processing. Recent work has explored its application for join discovery (Dong et al., 2022; Kayali et al., 2024), database tuning (Trummer, 2022), ML pipelines (Shankar and Parameswaran, 2024), natural language to SQL (Pourreza et al., 2024), and semantic table understanding (Fang et al., 2024; Cong et al., 2023). These efforts demonstrate the data management’s interest and success in applying LLMs (Fernandez et al., 2023b), but unlike DocETL, a general-purpose declarative system, they instead target specific problems with custom LLM implementations.
@@ 8. Conclusion
We introduced DocETL, a declarative system that optimizes complex document processing tasks using LLMs. By focusing on improving accuracy rather than just reducing costs, DocETL addresses critical limitations in existing LLM-powered data processing frameworks. We introduced several novel rewrite directives, an agent-based framework for plan rewriting and evaluation, and an opportunistic optimization strategy to handle the unique challenges of unstructured data analysis. Our evaluation across four unstructured document analysis tasks demonstrated that DocETL can find plans with outputs 25-80% more accurate than baselines. As LLMs continue to evolve and new challenges in complex document processing emerge, DocETL’s architecture provides a flexible foundation for future research and applications in complex document processing.

### v3

@@ 5.7. User Adoption and Impact
6. Discussion and Limitations
We reflect on the differences between DocETL and traditional database systems. Then, we acknowledge limitations of our work.
Revisiting Traditional Database Paradigms. LLM-powered document processing systems like DocETL differ from traditional database systems across several layers, ranging from operators—both logical and physical—to rewriting and optimization, to user specification and intent. The core difference is that LLMs are inherently non-deterministic and not always accurate. At the operator level, traditional database systems guarantee correct and consistent outputs, regardless of physical implementation choices. A hash join or nested loop join will produce identical results, differing only in performance characteristics. In contrast, DocETL’s output quality and correctness can vary dramatically based on physical operator parameter choices. For example, different fold batch sizes or gather configurations directly impact accuracy rather than just performance. Moreover, even logical operators work differently; unlike traditional DBMSes where atomic attributes (e.g., a blob of text) remain atomic, DocETL must decompose these attributes, because LLM accuracy often decreases as document size increases. The processing order of these decomposed units matters—LLMs cannot always process chunks independently, which motivates specialized operations like our gather operator and folding approach for reduce operations.
Second, rewriting in traditional databases is purely logical and guaranteed to produce identical output, but in DocETL, it’s semantic and instantiated by LLMs. In DocETL, operator output quality for rewrites can vary—e.g., if the LLM agent happens to synthesize bad prompts—and rewriting is not just for performance (latency), but also correctness. Moreover, unlike traditional DBMSes where there is only one way of executing a rewrite, here there are infinite. Consequently, the space of possible plans is infinite, even for a single operator, because each plan is instantiated by LLMs that could generate countless variations of prompts and configurations.
Third, optimization in DocETL is different because we lack a good model for accuracy, unlike well-understood cost models in traditional DBMS optimizers. DocETL must empirically evaluate plans using LLMs themselves as judges. LLMs are inherently uncalibrated for accuracy estimation (Liu et al., 2024c), so we must score and rank actual sample outputs using pairwise comparisons. Moreover, in DocETL, operator accuracy depends heavily on accuracy of other operators, and during optimization, the best suffix of a plan directly depends on which subplan was chosen for the prefix. This means that unlike traditional optimizers, e.g., Selinger’s algorithm (Selinger et al., 1979) or Cascades (Graefe, 1995), that can optimize subplans and apply dynamic programming or divide and conquer to assemble an optimal plan, DocETL cannot rely on operator independence.
Finally, the specification layer in DocETL is inherently fuzzy. DocETL specifications can be ambiguous because they are in natural language, allowing flexibility in interpretation by LLMs. However, this flexibility comes with limitations—specifications may not capture all corner cases that might arise during processing, requiring robust fault tolerance mechanisms to handle unexpected scenarios when output schemas fail or rate limits are exceeded.
Limitations. While DocETL demonstrates accuracy improvements in real-world applications, there are some limitations. First, using LLMs for both directive implementation and validation risks shared biases, although our validation agent’s different perspective often identifies issues despite sharing the underlying model. Recent work from the ML community has demonstrated that LLMs can often be more effective as verifiers than generators, especially if prompted to focus on specific criteria, as our validation agents do  (Zheng et al., 2023; Zhu et al., 2025; Zhang et al., [n. d.]). Nevertheless, our optimizer allows users to specify different models for the rewrite and validation agents, and we’re exploring several other mitigation strategies: diversifying prompts; incorporating human expertise in the optimizer through interactive refinement of pipelines; and exploring hybrid validation approaches that combine LLM assessment with traditional techniques (e.g., allowing a user to provide labeled ground truth to verify accuracy via exact string matching). These ideas are guiding our development of a prototype interface for DocETL that enables users to incorporate their expertise both during optimization and by directly editing or correcting outputs, providing a human-in-the-loop approach to complement automated complex document processing.
@@ 7. Related Work
LLM-powered data processing frameworks have recently gained significant attention in the database community. LOTUS (Patel et al., 2024) extends Pandas with semantic operators, while Palimpzest (Liu et al., 2024b) provides a declarative framework focusing on map-like operations. Aryn (Anderson et al., 2024) offers a Spark-like API with PDF extraction capabilities and human-in-the-loop processing. Unlike DocETL, these systems primarily make simplifying assumptions about task complexity, typically focusing on extraction tasks or queries that capable LLMs can handle without decomposition. They employ various cost-based optimizations, including classical techniques like predicate pushdown (Hellerstein and Stonebraker, 2005) and ML-specific approaches like model cascades (Wang et al., 2017). However, when applied to complex document processing tasks, even state-of-the-art models fall short. DocETL addresses this limitation through agent-driven optimization, exploring decomposition to improve accuracy. Moreover, DocETL is, uniquely, the only system to support documents with lengths that exceed LLM context windows and introduces new operators (e.g., gather, split) for this, as well as for entity resolution as a first-class citizen.
Other LLM-powered data processing systems focus on different settings—typically making strong assumptions about the structure of the documents and predictability of format. ZenDB (Lin et al., 2024) optimizes SQL queries for templatized documents, while DocETL handles arbitrary documents. EVAPORATE (Arora et al., 2023) specializes in table extraction through code synthesis (only where applicable, in semi-structured settings), which could complement DocETL. Regarding LLM agents: Caesura (Urban and Binnig, 2024) uses LLMs to translate natural language to SQL pipelines but leaves optimization for future work; CleanAgent (Qi and Wang, 2024) uses agents to standardize and clean data (and also does not consider optimization). Other systems propose specialized pipelines for specific tasks: for instance, Edge et al. (2024) use a fixed map-reduce pipeline with predefined prompts for knowledge graph querying—whereas DocETL enables flexible pipeline construction and optimization for any document processing task. A common limitation across these systems is inadequate context management, particularly for documents exceeding context windows or tasks requiring cross-document reasoning. While prompt optimization (Wen et al., 2024; Khattab et al., 2024) could complement DocETL, it falls short on complex document tasks, even with human guidance (White et al., 2023), particularly when data and tasks exceed single LLM call capabilities. Moreover, LLMs have been leveraged for a variety of data tasks beyond document processing, such as join discovery (Dong et al., 2022; Kayali et al., 2024), database tuning (Trummer, 2022), ML pipelines (Shankar and Parameswaran, 2024), natural language to SQL (Pourreza et al., 2024), semantic table understanding (Fang et al., 2024; Cong et al., 2023). and others (Fernandez et al., 2023b), but not for complex document processing.
@@ 8. Conclusion
We introduced DocETL, a declarative system that optimizes complex document processing tasks using LLMs. We introduced several novel rewrite directives, an agent-based framework for plan rewriting and evaluation, and an opportunistic optimization strategy. Our evaluation across four unstructured document analysis tasks demonstrated that DocETL can find plans with outputs 21-80% more accurate than baselines. DocETL is a first step toward an agentic optimizer for LLM-powered data processing. While exploring the large space of possible plan decompositions is hard, our approach shows that automated optimization is both feasible and beneficial. Future work will focus on reducing cost by considering cheaper models for simpler sub-tasks, and incorporating human feedback to refine plans. With growing real-world adoption, DocETL provides a foundation for future research and applications in complex document processing.
@@ A.2. Header Lineage Preservation
Appendix B Rewrite Directive Instantiation
This appendix provides a detailed walkthrough of how an LLM agent instantiates a rewrite directive in DocETL, using the legal contract analysis task in Section 5.1.
B.1. Initial Task and Baseline Operation
Our baseline approach used a single map operation with a prompt listing all 41 clause types to extract from legal contracts:
Given the following contract document:
{{ input.document }}
Extract the text spans (if they exist) for each of the following categories. If a category is not present or cannot be determined, return an empty string. If there are multiple text spans for a category, return them as a comma-separated list of text spans.
1. Document Name: The name of the contract
2. Parties: The two or more parties who signed the contract
3. Agreement Date: The date of the contract
4. Effective Date: The date when the contract is effective
5. Expiration Date: On what date will the contracts initial term expire?
6. Renewal Term: What is the renewal term after the initial term expires? This includes automatic extensions and unilateral extensions with prior notice.
7. Notice to Terminate Renewal: What is the notice period required to terminate renewal?
8. Governing Law: Which state/countrys law governs the interpretation of the contract?
9. Most Favored Nation: Is there a clause that if a third party gets better terms on the licensing or sale of technology/goods/services described in the contract, the buyer of such technology/goods/services under the contract shall be entitled to those better terms?
10. Non-Compete: Is there a restriction on the ability of a party to compete with the counterparty or operate in a certain geography or business or technology sector?
11. Exclusivity: Is there an exclusive dealing commitment with the counterparty? This includes a commitment to procure all "requirements" from one party of certain technology, goods, or services or a prohibition on licensing or selling technology, goods or services to third parties, or a prohibition on collaborating or working with other parties), whether during the contract or after the contract ends (or both).
12. No-Solicit of Customers: Is a party restricted from contracting or soliciting customers or partners of the counterparty, whether during the contract or after the contract ends (or both)?
13. Competitive Restriction Exception: This category includes the exceptions or carveouts to Non-Compete, Exclusivity and No-Solicit of Customers above.
14. No-Solicit of Employees: Is there a restriction on a partys soliciting or hiring employees and/or contractors from the counterparty, whether during the contract or after the contract ends (or both)?
15. Non-Disparagement: Is there a requirement on a party not to disparage the counterparty?
16. Termination for Convenience: Can a party terminate this contract without cause (solely by giving a notice and allowing a waiting period to expire)?
17. Right of First Refusal, Offer or Negotiation: Is there a clause granting one party a right of first refusal, right of first offer or right of first negotiation to purchase, license, market, or distribute equity interest, technology, assets, products or services?
18. Change of Control: Does one party have the right to terminate or is consent or notice required of the counterparty if such party undergoes a change of control, such as a merger, stock sale, transfer of all or substantially all of its assets or business, or assignment by operation of law?
19. Anti-Assignment: Is consent or notice required of a party if the contract is assigned to a third party?
20. Revenue/Profit Sharing: Is one party required to share revenue or profit with the counterparty for any technology, goods, or services?
21. Price Restriction: Is there a restriction on the ability of a party to raise or reduce prices of technology, goods, or services provided?
22. Minimum Commitment: Is there a minimum order size or minimum amount or units per-time period that one party must buy from the counterparty under the contract?
23. Volume Restriction: Is there a fee increase or consent requirement, etc. if one partys use of the product/services exceeds certain threshold?
24. IP Ownership Assignment: Does intellectual property created by one party become the property of the counterparty, either per the terms of the contract or upon the occurrence of certain events?
25. Joint IP Ownership: Is there any clause providing for joint or shared ownership of intellectual property between the parties to the contract?
26. License Grant: Does the contract contain a license granted by one party to its counterparty?
27. Non-Transferable License: Does the contract limit the ability of a party to transfer the license being granted to a third party?
28. Affiliate IP License-Licensor: Does the contract contain a license grant by affiliates of the licensor or that includes intellectual property of affiliates of the licensor?
29. Affiliate IP License-Licensee: Does the contract contain a license grant to a licensee (incl. sublicensor) and the affiliates of such licensee/sublicensor?
30. Unlimited/All-You-Can-Eat License: Is there a clause granting one party an "enterprise," "all you can eat" or unlimited usage license?
31. Irrevocable or Perpetual License: Does the contract contain a license grant that is irrevocable or perpetual?
32. Source Code Escrow: Is one party required to deposit its source code into escrow with a third party, which can be released to the counterparty upon the occurrence of certain events (bankruptcy, insolvency, etc.)?
33. Post-Termination Services: Is a party subject to obligations after the termination or expiration of a contract, including any post-termination transition, payment, transfer of IP, wind-down, last-buy, or similar commitments?
34. Audit Rights: Does a party have the right to audit the books, records, or physical locations of the counterparty to ensure compliance with the contract?
35. Uncapped Liability: Is a partys liability uncapped upon the breach of its obligation in the contract? This also includes uncap liability for a particular type of breach such as IP infringement or breach of confidentiality obligation.
36. Cap on Liability: Does the contract include a cap on liability upon the breach of a partys obligation? This includes time limitation for the counterparty to bring claims or maximum amount for recovery.
37. Liquidated Damages: Does the contract contain a clause that would award either party liquidated damages for breach or a fee upon the termination of a contract (termination fee)?
38. Warranty Duration: What is the duration of any warranty against defects or errors in technology, products, or services provided under the contract?
39. Insurance: Is there a requirement for insurance that must be maintained by one party for the benefit of the counterparty?
40. Covenant Not to Sue: Is a party restricted from contesting the validity of the counterpartys ownership of intellectual property or otherwise bringing a claim against the counterparty for matters unrelated to the contract?
41. Third Party Beneficiary: Is there a non-contracting party who is a beneficiary to some or all of the clauses in the contract and therefore can enforce its rights against a contracting party?
The output schema specified a list of clause-text pairs:
output:
schema:
clauses: list[{clause_type: string, text_span: string}]’
B.2. Optimization Process
B.2.1. Determine If Optimization Is Needed
The DocETL optimizer first evaluates whether the operation needed improvement. The system uses an LLM agent to synthesize a validator prompt. This agent is provided the operation prompt and sample outputs, and instructed to devise binary validation criteria around precision and recall. The following validator prompt was generated:
1. Did the operation successfully extract all relevant text spans for each
category listed in the task prompt, ensuring no instances were missed,
especially for categories with multiple possible text spans?
2. Is the extracted information presented in the correct format, with
text spans properly separated by commas where applicable, and are
there any inconsistencies or errors in the extracted data?
When applied to sample outputs from the original user-written operation, the LLM validator replies:
"Needs improvement. Row 1 output is mostly accurate, but it misses several important categories such as Renewal Term’, Notice to Terminate Renewal’, Governing Law’, Most Favored Nation’, Exclusivity’, No-Solicit of Customers’,... The extraction process should be improved to ensure that all categories listed in the task prompt are addressed, even if the information is not present in the document. For categories that do not have applicable text spans, the output should explicitly state that the category is not applicable or that the information is missing, rather than omitting it entirely."
B.2.2. Select and Apply Rewrite Directive
Based on this assessment, DocETL programmatically matches the case to applicable rewrite directives. For this example, we will focus on the implementation of Equation 11, which decomposes a complex map operation into independent subtasks. The optimizer then invokes the LLM agent with: the original prompt and output schema, sample contract documents, and instructions to decompose the task. As of March 2025,DocETL uses the following prompt to the LLM agent:
Original task prompt:
{original_prompt}
Output schema the operation will produce:
"clauses": "list[{clause_type: string, text_span: string}]"
Input data sample:
{sample_contract_text}
Decompose the original task into parallel subtasks, where each subtask
produces one or more keys of the output schema. Assume that the subtasks
can be executed independently. You cannot rely on the output of one subtask
to complete another subtask. Make sure you include the same input variables
as in the original task prompt. Each prompt should be a Jinja2 template...
[Additional instructions about format...]
B.2.3. LLM Agent Generates Decomposed Operations
The LLM agent responded with a detailed decomposition plan:
"subtasks": [
"name": "extract_basic_contract_information",
[…]
"output_keys": [
"document_name",
"parties",
"agreement_date",
"effective_date",
"expiration_date"
"name": "extract_renewal_and_termination_clauses",
[…]
"output_keys": [
"renewal_term",
"notice_to_terminate_renewal",
"termination_for_convenience"
"name": "extract_legal_governance_clauses",
[…]
"output_keys": [
"governing_law",
"anti_assignment",
"change_of_control"
....further clauses omitted due to brevity
We similarly prompt the LLM to synthesize an additional operation to convert the flat schema output from the parallel_map into the original output schema format expected by the user, following Equation 11. For the sake of this walk-through, we will assume that each newly synthesized operation is able to accurately perform its subtask, but in practice, DocETL will recursively apply rewrites to each new subtask.
B.2.4. Plan Evaluation and Selection
After generating the aforementioned plan and several alternatives (with variations in groupings and decomposition strategies), DocETL executes all the candidate plans on a sample, and invokes the validation agent to rate and compare outputs of plans. DocETL picks the top-ranked plan to replace the original operation with. Overall, the LLM agent effectively turns a high-level rewrite directive into a concrete implementation by, understanding the original task and data, identifying semantically meaningful ways to decompose it, generating appropriate prompts for each subtask, and creating the necessary schema transformations.